A brass tie bar of length 500 mm and diameter 20 mm undergoes a change in temperature from 25°C to 135°C. The linear expansivity of copper is 19 × 10−6°C−1 and its modulus of elasticity is 84 Gpa.


i)​Determine the change of length that takes place.

 

ii)​Determine the compressive force induced in the bar if it is rigidly held and expansion is prevented.



1
Expert's answer
2021-05-05T07:46:27-0400

i) dl = "\alpha"l∆T = 19 x 10-6 °C-1 x 500 mm x (135 - 25)°C = 1.045 mm

ii) F = "\sigma"A = "\alpha"TEA = 19 x 10-6 °C-1 x 84 x 109 Pa x (135 - 25)°C x "\pi"(0.01)2 = 55.15 kN